# Optimal Threshold for Fracton Codes and Nearly Saturated Code Capacity in Three Dimensions

### v1

@@ Abstract
Fracton codes have been intensively studied as novel topological states of matter, yet their fault-tolerant properties remain largely unexplored. Here, we investigate the optimal thresholds of self-dual fracton codes, in particular the checkerboard code, against stochastic Pauli noise. By utilizing a statistical-mechanical mapping combined with large-scale parallel tempering Monte Carlo simulations, we calculate the optimal code capacity of the checkerboard code to be $p_{th}\simeq 0.108(2)$. This value is the highest among known three-dimensional codes and nearly saturates the theoretical limit for topological codes. Our results further validate the generalized entropy relation for two mutually dual models, $H(p_{th})+H(\tilde{p}_{th})\approx 1$, and extend its applicability beyond standard topological codes. This verification indicates the Haah’s code also possesses a code capacity near the theoretical limit $p_{th}\approx 0.11$. These findings highlight fracton codes as highly resilient quantum memory and demonstrate the utility of duality techniques in analyzing intricate quantum error-correcting codes.
@@ I Introduction
Quantum error correction (QEC) plays a pivotal role in the development of large-scale quantum computing [Shor96, Knill97, Gottesman98, Dennis02]. Currently, the leading QEC strategy is the surface code, a variant of the two-dimensional toric code that has been realized across various quantum computing platforms [Google25, He25, Moses23, Bluvstein24]. Nevertheless, despite its structural simplicity and ease of experimental implementation, the surface code suffers from significant drawbacks, specifically regarding its inefficiency in encoding logical qubits and performing logical operations [Pantaleev22, Bravyi13, Pastawski15, Campbell17]. Consequently, the search for more efficient QEC schemes that surpass the performance of the surface code has become a major focus in the modern study of fault-tolerant quantum computation. Simultaneously, as the construction and fault-tolerance properties of error-correcting codes are deeply rooted in the physics of topological order and statistical mechanics [Dennis02, Bombin06, Bombin07a, Terhal15, Brown16], the exploration of new QEC strategies also holds promise for discovering novel phases and new physical phenomena.
@@
Over the past years, a wide array of quantum codes beyond the surface code has been proposed, including higher-dimensional variants of topological codes [Campbell17] and numerous quantum low-density parity-check codes utilizing remote interactions [Breuckmann21b]. Nevertheless, in stark contrast to the diversity of the codes, our understanding of their fault-tolerance properties remains rather limited. In particular, the construction primarily focuses on the code parameters $[[n,k,d]]$, where $n$ and $k$ represent the number of physical qubits and logical qubits, respectively, and $d$ denotes the code distance. While these parameters serve as key metrics for a code’s capability in encoding logical information, crucial metrics of the code’s resilience against errors are still lacking. Specifically, a code possessing favorable parameters may inherently suffer from a low error threshold, thereby rendering it practically irrelevant. Therefore, determining the fault-tolerance threshold is a critical subsequent step in characterizing these codes. Ideally, one would employ a decoding algorithm capable of handling realistic error models and identify the accurate threshold. However, due to the hypergraph structure inherent in the underlying decoding problem, efficient decoding algorithms remain unknown for the most of those codes. As demonstrated in various notable studies of topological codes, identifying the accurate threshold can prove highly non-trivial, even under the simplest error models.
@@
Fracton codes constitute a novel class of three-dimensional topological quantum codes that are of significant interest to both condensed matter physics and quantum computing [Haah11, Vijay16, Vijay15, Bravyi13_RG, Chamon05, Yoshida13, Ma17, Nandkishore19, Song19, WenDN, Aasen20, Pretko20, Zhou22, Song24]. In contrast to standard topological codes, which typically possess a constant ground-state degeneracy (GSD), fracton codes are characterized by a sub-extensive GSD and excitations that exhibit restricted mobility or immobility. Notably, previous studies on the X-cube fracton code have demonstrated strong error resilience, evidenced by a remarkably high threshold [Song22]. Specifically, the X-cube code achieves a code capacity threshold of approximately $0.075$, considerably surpassing the $0.033$ threshold of the $3D$ toric code [Wang03, Ohno2004] and the $0.017$ threshold of the $3D$ color code [Kubica18]. Nonetheless, although fracton codes have been extensively examined through the lenses of condensed matter theory, their fault-tolerant properties remain largely unexplored. To the best of our knowledge, the X-cube code represents the solitary case whose optimal threshold has been explicitly computed.
@@
   $2D$ Surface code    $0.1094(2)$ [Dennis02]    $0.1094(2)$ [Dennis02]    $0.9962(9)$
   $2D$ Color code    $0.109(2)$ [Katzgraber09]    $0.109(2)$ [Katzgraber09]    $0.994(9)$
   $3D$ Toric code    $0.2327(3)$ [Wang03]    $0.033(4)$ [Wang03]    $0.99(2)$
   $3D$ Color code    $0.276$ [Kubica18]    $0.019$ [Kubica18]    $0.986$
   X-Cube code    $0.152(4)$ [Song22]    $0.075(2)$ [Song22]    $1.00(1)$
   Checkerboard code    $\bf 0.108(2)$}    $\bf 0.108(2)$}    $\bf 0.987(8)$}
@@
Second, we corroborate the validity of a powerful generalized entropy duality across a broader scope of QEC codes. The quantum Gilbert-Varshamov (QGV) bound [Varshamov57, Calderbank96] indicates that the $X$ and $Z$ code capacities of a Calderbank–Shor–Steane (CSS) code satisfy the inequality $H(p_{th}^{X})+H(p_{th}^{Z})\leq 1$, where $H(p):=-p\log_{2}p-(1-p)\log_{2}(1-p)$ denotes the binary Shannon entropy. From the perspective of statistical physics, the saturation of this bound corresponds to the duality relation $H(p_{th})+H(\tilde{p}_{th})\approx 1$, where $p_{th}$ and $\tilde{p}_{th}$ denote the critical points of two mutually dual models [Nishimori07, Song22]. This implies that determining the critical point for one model allows for the deduction of the critical point for its dual. And for self-dual models, the critical point can be immediately estimated. Moreover, while this work focuses on code capacity, this duality extends to include scenarios involving measurement errors [Wang03, Xu25].
It is important to note that, although this generalized duality can save significant computational resources, it is rooted in a self-consistent replica-symmetry-breaking approximation [Nishimori07, Song22]. Therefore, it must be sufficiently verified before it can be confidently applied to different codes. Table 1 summarizes the success of this duality to standard topological codes. Combined with our current and previous findings, we confirm its applicability to fracton codes. Supported by these extensive justifications, we expect that Haah’s code, whose simulation is even more resource demanding, also nearly saturates the code capacity threshold $p_{th}\approx 0.11$.
@@ II Checkerboard code
The checkerboard code is defined on a cubic lattice of linear size $L$ with periodic boundary conditions (PBC), where a single qubit resides at each lattice site. The stabilizer operators $\hat{A}_{c}=\prod_{i\in\partial c}\hat{\sigma}_{i}^{x}$ and $\hat{B}_{c}=\prod_{i\in\partial c}\hat{\sigma}_{i}^{z}$ are defined at the center of every other cube $c$ as the tensor product of Pauli $\hat{\sigma}^{x}$ and $\hat{\sigma}^{z}$ operators between the 8 spins residing at its vertices. These stabilizers are distributed throughout the lattice in a checkerboard pattern as shown in Fig. 1(a), so that in a system of linear size $L$ there are $L^{3}/2$ $\hat{A}_{c}$ stabilizers and $L^{3}/2$ $\hat{B}_{c}$ stabilizers. To satisfy both periodic boundary conditions (PBC) and the checkerboard structure, the model is defined only for even $L$. These conditions ensure that neighbouring stabilizers always share 2 qubits, guaranteeing that all stabilizers commute with each other and that the model is exactly solvable, with the ground states satisfying $\hat{A}_{c}=\hat{B}_{c}=1,\forall c$.
@@
These fractons can be viewed as gapped topological defects generated on top of the ground state. Due to overlapping $\mathbb{Z}_{2}$ charge conservation laws, individual fractons cannot be moved without creating additional excitations. However, a pair of fractons (a dipole) possesses restricted mobility in directions orthogonal to its dipole moment. For example, as depicted in Fig. 1(b), an adjacent pair of fractons (a dipole) living in one $xz$-plane can be moved without energy cost along a rigid line $l$ in the $y$-direction by applying string operator $\prod_{n\in\textit{l}}\hat{\sigma}^{x}_{n}$ or $\prod_{n\in\textit{l}}\hat{\sigma}^{z}_{n}$. Moreover, if a dipole is oriented along the z-axis, the pair can move freely along the $x$ and $y$ directions. This motion is generated by stacks of adjacent $z$-strings, forming a flexible membrane that permits motion within the $xy$-plane, as shown in Fig. 1(c, d).
@@ III.1 Error Correction
For a given stabilizer model, let $Q$ denote the set of physical qubits. The commuting $X$-type and $Z$-type stabilizer operators are denoted by $\hat{A}_{\mathcal{S}_{X}}$ and $\hat{B}_{\mathcal{S}_{Z}}$, respectively, where $\mathcal{S}_{X/Z}$ serves as both a label for the stabilizer and an identifier for its anchor cell. The supports of these operators on the lattice are denoted by $\partial\mathcal{S}_{X},\partial\mathcal{S}_{Z}\subset Q$. The logical code space is defined as the ground-state subspace of the Hamiltonian:
@@
Error correction is asymptotically reliable if there exists a unique most-probable class $\eta_{s}^{*}]$ such that, in the thermodynamic limit, $\mathrm{Pr}([\eta_{s}^{*}])\to\mathrm{Pr}(s)$ [Dennis02]. This condition holds only when physical error rates remain below a threshold probability $p_{\text{th}}$, a limit guaranteed by the threshold theorem [Aharanov99]. Above this threshold, the logical error class cannot be reliably determined; thus, the optimal threshold is a fundamental parameter for assessing the performance of any quantum code.
@@ III.2 Statistical Mechanical Mapping
To ascertain whether it is possible to identify the most likely error class for a given physical error probability $p$, one can relate the quantum information properties of a given error model to the physical features of a classical Ising spin model subject to disorder. Ever since its first introduction in Ref. [Dennis02], this relation has been an active subject of study and has been formalized under the name of the statistical-mechanical (SM) mapping [Chubb18].
@@ III.3 Optimal Error Threshold
To determine the appropriate correction, it is necessary to identify the most probable equivalence class $\eta_{s}^{*}]_{X(Z)}$ such that, in the thermodynamic limit, $\Pr\left([\eta_{s}^{*}]_{X(Z)}\right)\to\Pr(s)$. For this to occur, the free energy difference $\delta F_{\eta,\lambda}$ associated with the addition of a logical operator must diverge in the thermodynamic limit. Since applying a logical operator corresponds to flipping a set of couplings across the entire lattice of the associated Ising model, the condition $\delta F_{\eta,\lambda}\to\infty$ implies that introducing a domain wall incurs a free energy cost proportional to its size. This condition is met only if the Hamiltonian $H_{\eta}$ supports an ordered phase at finite temperature [Dennis02]. The existence of an ordered phase at the temperature $\beta_{N}$ corresponding to the error probability $p$ ensures that, in the thermodynamic limit, the error model can always be decoded for error rates below the threshold. Thus, the optimal error threshold can be inferred by studying the order-disorder phase transition of the classical Ising model with quenched disorder.
@@ III.4 Generalized Duality in classical spin systems
While the disordered Ising model is characterized by the interplay of disorder and temperature, it accurately describes the statistics of the quantum spin system only when it satisfies the Nishimori condition $e^{-2\beta_{N}}=p/(1-p)$: this identifies the Nishimori line (NL), a special locus in parameter space along which the quenched disorder distribution is exactly matched to the Boltzmann weight of the classical model [Nishimori07]. Along the NL, the probability of flipping an Ising spin, dictated by its Boltzmann weight $e^{-\beta E}$, is identical to the probability of flipping the bonds which share said spin (which corresponds to changing the error configuration, dictated by the error probability $p$).
The enhanced symmetry between thermal and disorder fluctuations renders the Nishimori line a particularly interesting region of parameter space: beyond its already mentioned relation with the statistics of the corresponding QEC error model, this symmetry strongly suggests that the multicritical point of the disordered Ising model lies on the NL, owing to its invariance under renormalization-group transformations [Doussal89, Iba99]. This property implies the existence of a transition from an error correctable to a non-error correctable phase in the thermodynamic limit [Dennis02, Chubb18], which is guaranteed by the threshold theorem [Aharanov99]. On top of that, it is exactly along this line that it is possible to define an approximate fixed point condition which relates the position of the multicritical points of any Ising model with its geometrically dual model [Takeda05]. This can be derived by recalling the relation between the partition functions of the model and that of its dual.
@@
The geometrically dual model is constructed in the following way [Wegner71]: For each coupling $C_{j}$, define an Ising spin $\sigma_{j}^{*}$. Then, for each $\sigma_{i}$, introduce a coupling $C_{i}^{*}$ so that, for all $C_{j}$ with $\sigma_{i}\in\partial C_{j}$, we have $\sigma_{j}^{*}\in\partial C_{i}^{*}$.
With this lattice structure in mind, we can introduce for each $C_{j}^{*}$ the dual weights as the 2-component Fourier transform of the original weights $u^{*}_{\pm}(K)=\left(u_{+}(K)\pm u_{-}(K)\right)/\sqrt{2}$. Under this construction, the partition function of the original and dual lattice are related by the Normal Factor Graph (NFG) Duality theorem [Forney11] which, in this context, represents a generalized version of the Kramers-Wannier Duality [Kramers41a, Savit80]:
@@
where $a$ is some appropriate factor. It follows that the product $\mathbf{Z}\{u_{\pm}(K_{1})\}\tilde{\mathbf{Z}}\{u^{*}_{\pm}(K_{2})\}$ is invariant under the simultaneous transformations $u_{\pm}(K_{1})\rightleftharpoons u_{\pm}^{*}(K_{2})\,,\,\,u_{\pm}(K_{2})\rightleftharpoons u_{\pm}^{*}(K_{1})$: the transition points $K_{1},K_{2}$ of the two models are then given by the fixed point condition [Nishimori10]
@@
$=2^{-n/2}\left(1-p+(-1)^{k}p\right)\left(e^{K}+e^{-K}\right)^{n-k}\left(e^{K}-e^{-K}\right)^{k}.$
@@
$\frac{x_{k}^{*}}{x_{0}^{*}}=(1-2p)^{k}\,\,\text{(even)},\qquad\frac{x_{k}^{*}}{x_{0}^{*}}=(1-2p)^{k+1}\,\,\text{(odd)}.$ (14)
For $0<p<1$, the $k>0$ sectors are suppressed relative to the principal factor $x_{0}^{*}$ by at least a factor $(1-2p)^{2}<1$. Following the standard replica approach, we set $K=K_{N}(p)$ and impose the principal-factor condition only on the $k=0$ sector. This gives [Takeda05, Nishimori07, Song22]
@@
The inclusion of subleading sectors ($k>0$) would shift the estimate slightly while leaving the leading condition as the dominant constraint. Nevertheless, it was shown in several cases to yield the closest approximation of the true multicritical point [Hinczewski05, Hasenbusch07, Masayuki09, Kubica18, Gattringer18, Song22]. Explicitly, this results in
@@
where $H(p):=-p\,\text{log}p-(1-p)\,\text{log}(1-p)$. This corresponds to the limit of the QGV bound $H(p_{th}^{X})+H(p_{th}^{Z})\leq 1$. While the derivation of Eq. (16) relies on the leading Boltzmann factor, i.e., on the replica analysis, it works surprisingly well for known topological codes as summarized in Table 1.
With this in mind, we turn our attention to the $3D$ fracton models to determine whether the presence of subsystem symmetries may affect the reliability of this prediction. Specifically, as the non-self-dual X-cube code was shown to saturate the QGV bound in a previous work [Song22], here we complement it for the self-dual checkerboard code.

### v2

# arXiv:2512.22888v2[quant-ph] 28 Feb 2026

@@ Abstract
Fracton codes have been intensively studied as novel topological states of matter, yet their fault-tolerant properties remain largely unexplored. Here, we investigate the optimal thresholds of self-dual fracton codes, in particular the checkerboard code, against stochastic Pauli noise. By utilizing a statistical-mechanical mapping combined with large-scale parallel tempering Monte Carlo simulations, we calculate the optimal code-capacity threshold of the checkerboard code to be $p_{th}\simeq 0.107(3)$. This value is the highest among known three-dimensional codes and nearly saturates the theoretical limit for topological codes. Our results further validate the generalized entropy relation for two mutually dual models, $H(p_{th})+H(\tilde{p}_{th})\approx 1$, and extend its applicability beyond standard topological codes. This verification indicates the Haah’s code also possesses a code capacity threshold near the theoretical limit $p_{th}\approx 0.11$. These findings highlight fracton codes as highly resilient quantum memory and demonstrate the utility of duality techniques in analyzing intricate error-correcting codes.
@@ I Introduction
Quantum error correction (QEC) plays a pivotal role in the development of large-scale quantum computing [51, 34, 24, 20]. Currently, the leading QEC strategy is the surface code, a two-dimensional topological code that has been realized across various quantum computing platforms [2, 28, 40, 4]. Nevertheless, despite its structural simplicity and ease of experimental implementation, the surface code suffers from significant drawbacks, specifically regarding its inefficiency in encoding logical qubits and performing logical operations [47, 10, 48, 13]. Consequently, the search for more efficient QEC schemes that surpass the performance of the surface code has become a major focus in the modern study of fault-tolerant quantum computation. Simultaneously, as the construction and fault-tolerance properties of error-correcting codes are deeply rooted in the physics of topological order and statistical mechanics [20, 6, 7, 56, 12], the exploration of new QEC strategies also holds promise for discovering novel phases and new physical phenomena.
@@
Over the past years, a wide array of quantum codes beyond the surface code has been proposed, including higher-dimensional variants of topological codes [13] and numerous quantum low-density parity-check codes utilizing remote interactions [11]. Nevertheless, in stark contrast to the diversity of the codes, our understanding of their fault-tolerance properties remains rather limited. In particular, the construction primarily focuses on the code parameters $[[n,k,d]]$, where $n$ and $k$ represent the number of physical qubits and logical qubits, respectively, and $d$ denotes the code distance. While these parameters serve as key metrics for a code’s capability in encoding logical information, crucial metrics of the code’s resilience against errors are still lacking. Specifically, a code possessing favorable parameters may inherently suffer from a low error threshold, thereby rendering it practically irrelevant. Therefore, determining the fault-tolerance threshold is a critical subsequent step in characterizing these codes. Ideally, one would employ a decoding algorithm capable of handling realistic error models and identify the accurate threshold. However, due to the hypergraph structure inherent in the underlying decoding problem, efficient decoding algorithms remain unknown for the most of those codes. As demonstrated in various notable studies of topological codes [20, 44, 5, 36, 19, 53], identifying the accurate threshold can prove highly non-trivial, even under the simplest error models.
@@
Fracton codes constitute a novel class of three-dimensional topological quantum codes that are of significant interest to both condensed matter physics and quantum computing [25, 58, 57, 9, 17, 63, 39, 41, 52, 61, 1, 49, 64, 54]. In contrast to standard topological codes, which typically possess a constant ground-state degeneracy (GSD), fracton codes are characterized by a sub-extensive GSD and excitations that exhibit restricted mobility or immobility. Notably, previous studies on the X-cube fracton code have demonstrated strong error resilience, evidenced by a remarkably high threshold [53]. Specifically, the X-cube code achieves a code capacity threshold of approximately $0.075$, considerably surpassing the $0.033$ threshold of the $3D$ toric code [44] and the $0.019$ threshold of the $3D$ color code [36]. Nonetheless, although fracton codes have been extensively examined through the lenses of condensed matter theory, their fault-tolerant properties remain largely unexplored. To the best of our knowledge, the X-cube code represents the solitary case whose optimal threshold has been explicitly computed.
@@
   $2D$ Surface code    $0.1094(2)$ [20]    $0.1094(2)$ [20]    $0.9962(9)$
   $2D$ Color code    $0.109(2)$ [33]    $0.109(2)$ [33]    $0.994(9)$
   $3D$ Toric code    $0.2327(3)$ [46]    $0.033(4)$ [44]    $0.99(2)$
   $3D$ Color code    $0.276$ [36]    $0.019$ [36]    $0.986$
   X-Cube code    $0.152(4)$ [53]    $0.075(2)$ [53]    $1.00(1)$
   Checkerboard code    $\bf 0.107(3)$}    $\bf 0.107(3)$}    $\bf 0.98(2)$}
@@
Second, we corroborate the validity of a powerful generalized entropy duality across a broader scope of topological codes. It is expected that the $X$ and $Z$ code-capacity thresholds of a Calderbank–Shor–Steane (CSS) code with a zero rate satisfy the inequality $H(p_{th}^{X})+H(p_{th}^{Z})\leq 1$, where $H(p):=-p\log_{2}p-(1-p)\log_{2}(1-p)$ denotes the binary Shannon entropy [20, 23]. From the perspective of statistical physics, the saturation of this bound corresponds to the duality relation $H(p_{th})+H(\tilde{p}_{th})\approx 1$, where $p_{th}$ and $\tilde{p}_{th}$ denote the critical points of two mutually dual models [43, 53]. This implies that determining the critical point for one model allows for the deduction of the critical point for its dual. And for self-dual models, the critical point can be immediately estimated. Moreover, while this work focuses on code capacity thresholds, this duality extends to include scenarios involving measurement errors [59, 62].
It is important to note that, although this generalized duality can save significant computational resources, it is rooted in a self-consistent replica-symmetry-breaking approximation [43, 53]. Therefore, it must be sufficiently verified before it can be confidently applied to different codes. Table 1 summarizes the success of this duality to standard topological codes. Combined with our current and previous findings, we confirm its applicability to fracton codes. Supported by these extensive justifications, we expect that Haah’s code, whose simulation is even more resource demanding, also nearly saturates the code capacity threshold $p_{th}\approx 0.11$.
@@ II Checkerboard code
[…]
@@
These fractons can be viewed as gapped topological defects generated on top of the ground state. Due to overlapping $\mathbb{Z}_{2}$ charge conservation laws, individual fractons cannot be moved without creating additional excitations. However, a pair of fractons (a dipole) possesses restricted mobility in directions orthogonal to its dipole moment. For example, as depicted in Fig. 1(b), an adjacent pair of fractons (a dipole) living in one $xz$-plane can be moved without energy cost along a rigid line l in the $y$-direction by applying string operator $\prod_{n\in\textit{l}}\hat{\sigma}^{x}_{n}$ or $\prod_{n\in\textit{l}}\hat{\sigma}^{x/z}_{n}$. Moreover, if a dipole is oriented along the z-axis, the pair can move freely along the $x$ and $y$ directions. This motion is generated by stacks of adjacent $z$-strings, forming a flexible membrane that permits motion within the $xy$-plane, as shown in Fig. 1(c, d).
@@ III.1 Error Correction
For a given stabilizer model, let $Q$ denote the set of physical qubits. The commuting $X$-type and $Z$-type stabilizer operators are denoted by $\hat{A}_{\mathcal{S}_{X}}$ and $\hat{B}_{\mathcal{S}_{Z}}$, respectively, where $\mathcal{S}_{X/Z}$ serves as both a label for the stabilizer and an identifier for its anchor cell (in the Checkerboard model, these would be the cubic cells $c$ as shown in Fig. 1). The supports of these operators on the lattice are denoted by $\partial\mathcal{S}_{X},\partial\mathcal{S}_{Z}\subset Q$. The logical code space is defined as the ground-state subspace of the Hamiltonian:
@@
Error correction is asymptotically reliable if there exists a unique most-probable class $\eta_{s}^{*}]$ such that, in the thermodynamic limit, $\mathrm{Pr}([\eta_{s}^{*}])\to\mathrm{Pr}(s)$ [20]. This condition holds only when physical error rates remain below a threshold probability $p_{\text{th}}$, a limit guaranteed by the threshold theorem [3]. Above this threshold, the logical error class cannot be reliably determined; thus, the optimal threshold is a fundamental parameter for assessing the performance of any quantum code.
@@ III.2 Statistical Mechanical Mapping
To ascertain whether it is possible to identify the most likely error class for a given physical error probability $p$, one can relate the quantum information properties of a given error model to the physical features of a classical Ising spin model subject to disorder. Ever since its first introduction in Ref. [20], this relation has been an active subject of study and has been formalized under the name of the statistical-mechanical (SM) mapping [19].
@@ III.3 Optimal Error Threshold
To determine the appropriate correction, it is necessary to identify the most probable equivalence class $\eta_{s}^{*}]_{X(Z)}$ such that, in the thermodynamic limit, $\Pr\left([\eta_{s}^{*}]_{X(Z)}\right)\to\Pr(s)$. For this to occur, the free energy difference $\delta F_{\eta,\lambda}$ associated with the addition of a logical operator must diverge in the thermodynamic limit. Since applying a logical operator corresponds to flipping a set of couplings across the entire lattice of the associated Ising model, the condition $\delta F_{\eta,\lambda}\to\infty$ implies that introducing a domain wall incurs a free energy cost proportional to its size. This condition is met only if the Hamiltonian $H_{\eta}$ supports an ordered phase at finite temperature [20]. The existence of an ordered phase at the temperature $\beta_{N}$ corresponding to the error probability $p$ ensures that, in the thermodynamic limit, the error model can always be decoded for error rates below the threshold. Thus, the optimal error threshold can be inferred by studying the order-disorder phase transition of the classical Ising model with quenched disorder.
@@ III.4 Generalized Duality in classical spin systems
While the disordered Ising model is characterized by the interplay of disorder and temperature, it accurately describes the statistics of the quantum spin system only when it satisfies the Nishimori condition $e^{-2\beta_{N}}=p/(1-p)$: this identifies the Nishimori line (NL), a special locus in parameter space along which the quenched disorder distribution is exactly matched to the Boltzmann weight of the classical model [43]. Along the NL, the probability of flipping an Ising spin, dictated by its Boltzmann weight $e^{-\beta E}$, is identical to the probability of flipping the bonds which share said spin (which corresponds to changing the error configuration, dictated by the error probability $p$).
The enhanced symmetry between thermal and disorder fluctuations renders the Nishimori line a particularly interesting region of parameter space: beyond its already mentioned relation with the statistics of the corresponding QEC error model, this symmetry strongly suggests that the multicritical point of the disordered Ising model lies on the NL, owing to its invariance under renormalization-group transformations [37, 30]. This property implies the existence of a transition from an error correctable to a non-error correctable phase in the thermodynamic limit [20, 19], which is guaranteed by the threshold theorem [3]. On top of that, it is exactly along this line that it is possible to define an approximate fixed point condition which relates the position of the multicritical points of any Ising model with its geometrically dual model [55]. This can be derived by recalling the relation between the partition functions of the model and that of its dual.
@@
The geometrically dual model is constructed using the procedure outlined in [60] and consists of a set of dual Ising spins $\sigma^{*}$ connected by a set of dual couplings $C^{*}$. We can introduce the dual weights associated to each coupling of the dual model as the 2-component Fourier transform of the original weights $u^{*}_{\pm}(K)=\left(u_{+}(K)\pm u_{-}(K)\right)/\sqrt{2}$. Under this construction, the partition function of the original and dual lattice are related by the Normal Factor Graph (NFG) Duality theorem [21] which, in this context, represents a generalized version of the Kramers-Wannier Duality [35, 50]:
@@
where $a$ is some appropriate factor. It follows that the product $\mathbf{Z}\{u_{\pm}(K_{1})\}\tilde{\mathbf{Z}}\{u^{*}_{\pm}(K_{2})\}$ is invariant under the simultaneous transformations $u_{\pm}(K_{1})\rightleftharpoons u_{\pm}^{*}(K_{2})\,,\,\,u_{\pm}(K_{2})\rightleftharpoons u_{\pm}^{*}(K_{1})$: the transition points $K_{1},K_{2}$ of the two models are then given by the fixed point condition [42]
@@
$=2^{-n/2}\left(p+(-1)^{k}(1-p)\right)\left(e^{K}+e^{-K}\right)^{n-k}\left(e^{K}-e^{-K}\right)^{k}.$
@@
$\frac{x_{k}^{*}}{x_{0}^{*}}=(1-2p)^{k}\,\,\text{(even)},\qquad\frac{x_{k}^{*}}{x_{0}^{*}}=-(1-2p)^{k+1}\,\,\text{(odd)}.$ (14)
For $0<p<1$, the $k>0$ sectors are suppressed in magnitude relative to the principal factor, with $|x_{k}^{*}/x_{0}^{*}|\leq(1-2p)^{2}<1$. Following the standard replica approach, we set $K=K_{N}(p)$ and impose the principal-factor condition only on the $k=0$ sector. This gives [55, 43, 53]
@@
The inclusion of subleading sectors ($k>0$) would shift the estimate slightly while leaving the leading condition as the dominant constraint. Nevertheless, it was shown in several cases to yield the closest approximation of the true multicritical point [29, 27, 45, 36, 22, 53]. Employing the replica trick limit $n\rightarrow 0$, the principal factor condition reduces to
@@
where $H(p):=-p\,\text{log}p-(1-p)\,\text{log}(1-p)$. This corresponds to the limit of the bound $H(p_{th}^{X})+H(p_{th}^{Z})\leq 1$. While the derivation of Eq. (16) relies on the leading Boltzmann factor, i.e., on the replica analysis, it works surprisingly well for known topological codes as summarized in Table 1.
With this in mind, we turn our attention to the $3D$ fracton models to determine whether the presence of subsystem symmetries may affect the reliability of this prediction. Specifically, as the non-self-dual X-cube code was shown to saturate the error threshold bound in a previous work [53], here we complement it for the self-dual checkerboard code.